OpenStudy (anonymous):

2x+2y=6 -3x+5y=-33 Linear System by using Substitution

OpenStudy (saifoo.khan):

x = 6 y = -3

OpenStudy (anonymous):

can u plz help me with the steps I am kinda a visual learner.

OpenStudy (anonymous):

Multiply first equation by 3 and second equation by 2

OpenStudy (anonymous):

When you add, you would have cancel out the x, only leaving the y

OpenStudy (anonymous):

That's should get you started

OpenStudy (anonymous):

time (3) 2x+2y=6 6x +6y = 18 time (2) -3x+5y=-33 -6x +10y = -66 Add ---------- 16y = -48 y = -48/16 y= -3

OpenStudy (anonymous):

substitution y=-3 find x 2x+2y=6 2x+2(-3)=6 2x-6=6 2x=6+6 x=6

OpenStudy (anonymous):

so how did you know to mutiply by the 3&the 2

OpenStudy (anonymous):

1) Solve either equation in terms of either x or y. I'll solve 2x+2y=6 in terms of x. You get: x = -y + 3 2) Substitute this back into the other equation -3(-y+3)+5y=-33 3) Solve for y y = -3 4) Since y = -3 and we know that x = -y + 3.... x = 6

OpenStudy (anonymous):

nancy can u plz explain to me how you knew to multiply by the 3 & 2=)